OpenStudy (anonymous):

Let f(x) = x – 3 and g(x) = 2x + 4. What is f(g(x)) ?

OpenStudy (anonymous):

I think I remember how to do this... f(g(x)) = -3 (2x + 4) f(g(x) = -6x -12

OpenStudy (anonymous):

f(g(x)) = (2x+4)-3 = 2x+1

OpenStudy (anonymous):

f(x) = x - 3 f(4) means plug in 4 everywhere there is an X f(g(x)) means plug in g(x) everywhere you see an X since g(x) = 2x + 4 f(g(x)) = (2x+4) - 3